# Gradient-free neural topology optimization

### v1

@@ Abstract
Gradient-free optimizers allow for tackling problems regardless of the smoothness or differentiability of their objective function, but they require many more iterations to converge when compared to gradient-based algorithms. This has made them unviable for topology optimization due to the high computational cost per iteration and high dimensionality of these problems. We propose a pre-trained neural reparameterization strategy that leads to at least one order of magnitude decrease in iteration count when optimizing the designs in latent space, as opposed to the conventional approach without latent reparameterization. We demonstrate this via extensive computational experiments in- and out-of-distribution with the training data. Although gradient-based topology optimization is still more efficient for differentiable problems, such as compliance optimization of structures, we believe this work will open up a new path for problems where gradient information is not readily available (e.g. fracture).
Keywords Gradient-free optimization  $\cdot$ Generative deep learning  $\cdot$ Latent optimization  $\cdot$ Topology Optimization
@@ 1 Introduction
Topology optimization plays an important role in the design of structures. Historically, method development has been focused on relatively simple problems, such as compliance minimization that are linear and differentiable [1, 2, 3]. While these problems are of high importance, many practical scenarios are governed by non-differentiable, non-linear objectives that are much more challenging to optimize for [4, 5, 6]. In principle, gradient-free optimization algorithms could be a promising approach for solving such problems because they can be applied to any objective. However, they suffer from a major limitation: they require several orders of magnitude more objective evaluations (simulations) than gradient-based optimizers to converge to a solution.
@@
Gradient-free optimizers update the solution by sampling and comparing the performance of trial solutions. At each iteration, the objective needs to be evaluated for the whole population of samples, using expensive simulations (usually FEM). Moreover, gradient-free optimizers suffer from the curse of dimensionality, i.e. the cost of optimization grows exponentially with the number of design variables [7, 8]. This is especially nefarious to topology optimization, as the number of design variables is typically very large rendering these optimizers unfeasible [1, 8].
@@
Considering these limitations, it comes as no surprise that gradient-free optimization is not well motivated for most topology optimization problems [8]. However, there are scenarios where the use of gradient-based methods might be challenging, namely in cases where the gradients are difficult to compute or where the objective function is noisy or discontinuous [4, 9]. In principle, gradient-free methods would be a logic solution in these contexts, but their slow convergence detracts from their use [8].
Contributions
@@
In this article we address the large rift in performance between gradient-free and gradient-based topology optimization for problems where the latter is orders of magnitude more efficient than the former. Therefore, we consider thousands of topology design problems where gradient-based optimizers severely outperform the best gradient-free algorithms and show that an appropriate machine learning strategy combining training (offline) and neural reparameterization (online) closes the performance gap by an order of magnitude.
@@
Compared to the state-of-the-art latent optimization of topology with gradient-free methods [10], our work explicitly targets the issue of scalability of gradient-free optimization. We show that the appropriate choice of the model architecture leads to significant gains in performance, compared to the conventional approach (without latent space reparameterization), but also compared to more standard architectures.
@@
We test our approach on out-of-distribution examples, i.e. examples that were not used in training and that include characteristics that were not explored in training. Furthermore, we deploy our model on an optimization problem with different physics and significantly different design features, to demonstrate promising generalization performance. Notwithstanding, we do not claim that the proposed approach can outperform gradient-based methods in convex or nearly convex objective functions, as occurs in the compliance optimization problems that we consider in this work. However, we believe that this research is a first step towards solving more challenging problems in the future, involving plasticity, fracture, and fatigue.
@@ 2 Related work
2.1 Limitations of gradient-based approaches
In many optimization problems the objective function is differentiable, such as optimization of compliance, and the derivatives can be obtained analytically or with automatic differentiation [11, 12, 1, 13]. In case the function is non-differentiable, so far, the most widespread approach is to linearize the problem such that it can still be solved with a gradient-based optimizer, in particular using the adjoint method [14, 6, 5, 15, 16]. Although relatively successful, this approach comes with certain limitations, as highlighted for example in [4], where the authors use the adjoint formulation of the topology optimization problem to postpone fracture. The presented derivation of the adjoint assumes that the objective function is smooth, however, this is not the case for fracture, where even the smallest changes in the design can have a critical impact on the performance (specifically, these changes can completely change the crack trajectory). As highlighted by the authors, the effects cannot be easily remedied by fixing the step size of the descent algorithm or by regularization. In such a case, when the objective function is not well behaved, discontinuous, with multiple local minima, and challenging to regularize, gradient-free methods could be a promising alternative [8, 4].
Several gradient-free algorithms were explored in topology optimization, including Genetic Algorithms [17, 18], Artificial Immune Algorithms [19], Ant Colonies [20, 21], Particle Swarms [22], Simulated Annealing [23], Harmony Search [24], Differential Evolution [25] among others [8]. These approaches, however, received strong criticism in the community [8] primarily due to their inferior efficiency, as opposed to gradient-based methods. Gradient-free algorithms suffer from the curse of dimensionality [8, 7], which means that the cost of optimization increases approximately exponentially with the number of problem dimensions. In standard topology optimization approaches [11], where the design variables correspond to element-wise densities, the dimensionality of even small ’toy-examples’ is in practice too large for gradient-free algorithms, leading to prohibitive costs (in order of 10,000 - 100,000 objective function evaluations, as opposed to the order of 10 – 100 evaluations for gradient-based methods [8]).
@@ 2.2 Latent-space optimization
The curse of dimensionality of gradient-free optimizers can be addressed by reparameterizing the optimization problem into a lower-dimensional latent space [26, 27, 28]. This dimensionality reduction can be well accommodated using generative deep learning models, such as generative adversarial networks (GAN) [29], variational autoencoders [30], and diffusion models [31, 32]. In the context of latent space optimization, variational autoencoders are a particularly common choice [33, 34].
Optimizing in latent space of a VAE was demonstrated in context of topology optimization [35, 36, 10]. Sato et al [35] and Gladstone et al. [36] explored latent optimization with gradient-based approach using a surrogate model – a neural network trained to predict the property of interest from the latent representation. The method takes advantage of the differentiability of the surrogate - as the property of interest can be minimized by backpropagating through the surrogate. In both cases, the results demonstrate that VAE is capable of representing new designs, that outperform those seen in the training. The disadvantage of this approach, however, is the need for data, required for training the surrogate.
Guo et al. [10] explore topology optimization in the latent space of a VAE with different optimization algorithms, including a gradient-free genetic algorithm. Unlike other works, the training of the VAE was augmented with an additional style-transfer procedure, to improve the quality of the generated designs. The training dataset comprised designs optimized for thermal compliance using the SIMP method [11], but the model was then used to solve a different problem - a multiobjective optimization, minimizing the maximum temperature and maximizing the power density. Similar to other works [35, 36], the method was able to synthesize novel designs, different to those seen in the training dataset. Most importantly, the results demonstrate that the pre-trained model can be successfully deployed on a different optimization problem, given the designs have similar features. Nevertheless, in the presented experiments the optimization process with the gradient-free method (Genetic Algorithm) required an order of 30,000 FEM evaluations, indicating relatively limited improvement as compared to the gradient-free optimization without reparameterization into the latent space [8].
@@ 2.3 Machine learning in topology optimization
In the context of topology optimization, generative machine learning models have been primarily used to map boundary conditions to final designs [37, 2, 38]. Different generative models were explored in such a setup, including GANs [39, 40] and diffusion models [41]. While demonstrating competitive performance, these approaches are often criticized for their limited generalization [37, 2], as they remain applicable only to the problem where they were trained on. In order to apply such a model to a new type of design problem (e.g. different objective, different physics), the model needs to be retrained on a dataset corresponding to the new problem. Generating the dataset, however, requires the use of conventional optimization methods and simulations.
Another relevant stream of work is neural reparameterization [12, 42, 13, 43, 2], where the neural network is used to implicitly represent the designed structure. Instead of optimizing the element-wise density values, the optimizer adjusts the trainable parameters of a neural network which outputs the density field. This approach proved beneficial in gradient-based approaches, where the weights can be adjusted by back-propagating the gradients of the density field (design sensitivities) [12]. Neural reparameterization was not explored in a gradient-free optimization scenario. One reason is the relatively high number of parameters of the neural networks, ranging from 100s [43] to 100,000s [44], making most network architectures unfeasible to optimize with gradient-free algorithms.
@@ 3 Methods
We propose a latent space optimization approach following two steps, as shown on the schematic in Figure 1. First, we train a generative model to reparameterize the topology designs into a lower-dimensional latent representation. Then, in the second step, the latent space is explored by the gradient-free optimizer, which attempts to find a design minimizing the objective. The optimizer controls the latent vector which parameterizes the design. This vector is decoded into a physical design using the pre-trained generative model, and the design is evaluated using the objective function (e.g. compliance). During one iteration of optimization the optimizer generates a population of trial latent vectors, which are subsequently evaluated, and based on their associated objective values, the optimizer updates the population according to its internal update rules, converging to a better solution.
(a) Schematic of a variational autoencoder (VAE). In the first step of our method we train the generative model - a VAE to reparameterize topology designs into latent space. Here $\mu$ and $\sigma$ are mean and variance vectors in latent space, $\epsilon$ is the noise vector sampled from a multivariate Gaussian, and $z$ is a latent variable.
(b) Schematic of latent space optimization process with gradient-free optimizer. In the second step of our method, we use a trained generative model, to optimize the designs using latent space representation of the generative model. Note that the latent vector $z$ is no longer stochastic, and it is fully controlled and updated ($z_{new}$) by the optimizer, based on the objective values obtained from the simulator.
Figure 1: Schematics of a) the generative model and b) the latent space optimization setup used in our proposed methods.
The key assumption of our approach is that the generative model learns compact representations of features occurring in structural designs, but these designs, in principle, are not necessarily bound to any particular objective function. We observe that the designs optimized for different objectives (e.g. compliance, minimum mass, or compliant mechanism design) oftentimes share the same set of visual features (e.g. compare the designs optimized for fracture to those optimized for standard compliance in [4]). This observation can be verified using quantitative feature similarity metrics such as Learned Perceptual Image Patch similarity [45] (for details, see the appendix). Based on that, we assume that the generator trained to capture visual features can be deployed to optimize different objectives without the need for retraining, and is not strictly limited to the problems used for generating the training dataset. This is a crucial advantage over state-of-the-art approaches exploring generative models for topology optimization.
@@ 3.1 Generative Model
We propose the use of a new VAE variant called Latent Bernoulli Autoencoder (LBAE) because it uses the Bernoulli distribution for modeling the latent space [46]. Although this architecture was proposed in a different context, it has been shown to achieve superior performance compared to conventional VAE architectures in terms of quality of the generated samples (FID score). We conjectured that using a Bernoulli distribution for the latent space (zeros and ones), instead of a continuous distribution such as a Gaussian, is particularly adequate for the case of topology optimization where the aim is to decide whether each discrete element should contain material or not. We carefully compare the LBAE to a conventional VAE architecture and searched for the best hyperparameters for a large class of problems, in an attempt to provide consider each method at their best performance [10]. Furthermore, we include gradient-based results and different gradient-free optimizers to establish clear performance metrics and clarify the usefulness of the proposed method.
@@
Conventionally, VAE provides a generative model that captures the distribution of data $\mathcal{D}=\left\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},...,\mathbf{x}^{(N)}\right\}$, by assuming a joint distribution $p(\mathbf{x},\mathbf{z})$ using some hidden (latent) variable $\mathbf{z}$, such that given $\mathbf{z}$, it can generate samples of $\mathbf{x}$. In practice, this is achieved by using two neural networks: the encoder, which approximates the posterior $p(\mathbf{z}|\mathbf{x})\approx q_{\phi}(\mathbf{z}|\mathbf{x})$, (since in general $p(\mathbf{z}|\mathbf{x})$ is intractable), and the decoder, which captures $p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})$. The two networks are trained simultaneously by maximizing the evidence lower bound, as shown in Equation 1.
$L_{\mathbf{\theta},\mathbf{\phi}}=\operatorname{\mathbb{E}}_{q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\frac{p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})p(\mathbf{z})}{q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})}\right]=\operatorname{\mathbb{E}}_{q(\mathbf{z}|\mathbf{x})}\left[\log p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})\right]-D_{KL}\left(q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})\right)$ (1)
@@
The first term corresponds to the reconstruction loss, and the second term corresponds to the Kullback-Leibler divergence, which regularizes the latent variable and aims to drive the parameterized distribution towards the assumed prior. In a conventional VAE, the prior $p(\mathbf{z})$ is assumed to be a multivariate Gaussian, while in LBAE the latent space is assumed to be a multivariate Bernoulli distribution. Unlike in VAE, however, in LBAE the Bernoulli distribution of the latent variable is imposed via network architecture and its specific sampling routine, rather than by an explicit regularization term (for details see [46]). Therefore, the second term from Equation 1 is not present in the LBAE loss.
@@
The training dataset $x\sim X$ in our case corresponds to images representing different topology designs. Once trained, the autoencoder is then used for generating new designs by sampling $z$ from the latent space. In this way, the designs can be represented using a lower dimensional latent vector.
@@
The LBAE architecture that we use is based on the model proposed by Fajtl et al. [46], adapted to our dataset by adjusting the number of layers and convolutional filters. According to the authors [46], modeling latent space with Bernoulli distribution produces sharper images, compared to traditional VAE, and allows for smoother interpolation in the latent space. We hypothesize that these properties are crucial for optimization, allowing for easier traversing of the loss landscape by the optimizer. Furthermore, unlike in the baseline VAE, LBAE uses residual layers. Combined with the Bernoulli distribution of the latent variable, we observed that this architecture indeed is capable of producing outputs with finer detail and reduced blurriness, addressing one of the main shortcomings of VAE [2]. The baseline VAE architecture used in our case is a modified model of Larsen et al. [47]. The inputs and outputs are single-channel, grayscale representations of the designs of resolution 64x64.
For both LBAE and VAE, one of the most important parameters is the dimensionality of the latent space, which balances the training and the optimization performance. While increasing the latent space dimensionality makes the training easier, as the compression of information is not as severe, the optimization becomes less efficient due to the curse of dimensionality [8, 7]. To find a balanced latent space dimensionality, we tested several configurations (for details, see the appendix). In the case of VAE, we did not observe a significant difference between the dimensionality 32 and 64, instead, we noticed that the performance was limited by insufficient resolution of finer details in the output (caused by the blurriness intrinsic to the VAE). These limiting factors, however, were to a large extent alleviated in LBAE, where the trade-off due to the latent space dimensionality was much more clear (for details, see the appendix). We found that for LBAE the dimensionality 256 offered the best balance.
@@
We use two different datasets of topology designs at train and test time, respectively. The training dataset was generated based on a modified approach of [48] and [41], resulting in 48,872 designs. The boundary conditions were generated as randomly drawn point loads and point supports, using Poisson’s distribution specified in the same way as in [48]. Additionally, in our dataset, we randomly applied a square mask to restrict part of the design domain from applying material (e.g. for an L-shaped beam the mask would be applied in the corner). The mask was applied with 25% probability, and the position of the mask was randomly assigned along one of the edges of the design domain. Volume fractions were drawn using a uniform distribution with a range of [0.12, 0.5]. This dataset was split into two subsets: for training and validation, using a standard 80:20 split.
The testing dataset was used only for monitoring the testing loss, which aims to verify the generalization capability of the model. This dataset was generated using the problems defined in [12]. The boundary conditions were generated using one of 28 classes of parameterized problems, with the parameters drawn at random from a uniform distribution, which resulted in over 4000 instances of testing problems. Unlike the training problem set, the testing problem classes include distributed loads and supports, therefore they are expected to provide a good insight into the generalization of the model and can be considered out-of-distribution.
The designs in both datasets were obtained using a gradient-based MMA optimizer with the default hyperparameter settings. The final designs were represented as grayscale images of fixed resolution (64x64) with pixel values representing the volume fraction of the material. For each problem, the optimization with MMA [49] was run for 200 iterations. The simulations were set up following the approach of the 88-line code [11], with $E_{0}=1.0$, $E_{min}=1e-9$, $\nu=0.3$ and a standard cone filter with a radius of 2 pixels for regularizing the density field.

### v2

@@ Abstract
Gradient-free optimizers allow for tackling problems regardless of the smoothness or differentiability of their objective function, but they require many more iterations to converge when compared to gradient-based algorithms. This has made them unviable for topology optimization due to the high computational cost per iteration and the high dimensionality of these problems. We propose a gradient-free neural topology optimization method using a pre-trained neural reparameterization strategy that addresses two key challenges in the literature. First, the method leads to at least one order of magnitude decrease in iteration count to reach minimum compliance when optimizing designs in latent space, as opposed to the conventional gradient-free approach without latent parameterization. This helps to bridge the large performance gap between gradient-free and gradient-based topology optimization for smooth and differentiable problems like compliance optimization, as demonstrated via extensive computational experiments in- and out-of-distribution with the training data. Second, we also show that the proposed method can optimize toughness of a structure undergoing brittle fracture more effectively than a traditional gradient-based optimizer, delivering an objective improvement in the order of 30% for all tested configurations. Although gradient-based topology optimization is more efficient for problems that are differentiable and well-behaved, such as compliance optimization, we believe that this work opens up a new path for problems where gradient-based algorithms have limitations.
Keywords Gradient-free optimization  $\cdot$ Generative deep learning  $\cdot$ Latent optimization  $\cdot$ Compliance  $\cdot$ Fracture  $\cdot$ Fatigue  $\cdot$ Plasticity
@@ 1 Introduction
Topology optimization plays an important role in the design of structures [1, 2, 3]. Historically, interest in gradient-free optimization has been present since the early days of the topology optimization field [4, 5], motivated mostly by their ability to deal with discrete design variables and their generality. However, with the advancements of the continuous design variable formulation, the developments focused on linear and differentiable problems, such as compliance minimization – suitable for a more efficient gradient-based optimization [1, 6, 7, 8]. While these problems are of high academic importance, many practical scenarios are governed by non-differentiable, non-linear objectives that are much more challenging to optimize for [9, 10, 11, 12]. In principle, gradient-free optimization algorithms could be a promising approach for solving such problems because they can be applied to any objective, thus remaining a relevant alternative. However, they suffer from a major limitation: they require several orders of magnitude more objective evaluations (simulations) than gradient-based optimizers to converge to a solution [13, 12].
@@
Gradient-free optimizers update the solution by sampling and comparing the performance of trial solutions. At each iteration, the objective needs to be evaluated for the whole population of samples, using expensive simulations (usually FEM). Moreover, gradient-free optimizers suffer from the curse of dimensionality, i.e. the cost of optimization grows exponentially with the number of design variables [14, 13]. This is especially nefarious to topology optimization, as the number of design variables is typically very large, rendering these optimizers unfeasible [1, 13].
@@
Taking into account these limitations, it comes as no surprise that gradient-free optimization is not well motivated for most topology optimization problems [13]. However, there are scenarios where the use of gradient-based methods might be challenging, namely in cases where the gradients are difficult to compute or where the objective function is noisy or discontinuous [9, 15, 12, 16]. In principle, gradient-free methods would be a logical solution in these contexts, but their slow convergence detracts from their use [13, 12].
@@
In this article, we address the large rift in performance between gradient-free and gradient-based topology optimization for problems where the latter is orders of magnitude more efficient than the former. Therefore, we consider thousands of topology design problems where gradient-based optimizers severely outperform the best gradient-free algorithms and show that an appropriate machine learning strategy combining training (offline) and neural reparameterization (online) closes the performance gap by an order of magnitude.
@@
Compared to the state-of-the-art latent optimization of topology with gradient-free methods [17], our work explicitly targets the issue of scalability of gradient-free optimization. We show that the appropriate choice of the model architecture leads to significant gains in performance, compared to the conventional approach (without latent space reparameterization), but also compared to more standard architectures. We test our approach on out-of-distribution examples, i.e. examples that were not used in training and that include characteristics that were not explored in training.
@@
We demonstrate the effectiveness of our proposed approach by deploying it on a non-linear path-dependent optimization problem that proves to be challenging for standard gradient-based approaches – namely optimizing structures undergoing brittle fracture. We test different initialization strategies and show that the gradient-based optimization remains prone to getting stuck in local minima, while the gradient-free optimizer coupled with latent reparameterization delivers improved performance without retraining or tuning the hyperparameters. Notwithstanding, we do not claim that the proposed approach can outperform gradient-based methods in convex or nearly convex objective functions, as occurs in compliance optimization problems. However, we believe that this research is a first step towards solving more challenging problems in the future, involving plasticity, fracture, and fatigue.
@@ 2 Related work
2.1 Limitations of gradient-based and gradient-free approaches
In many optimization problems, the objective function is differentiable, such as optimization of compliance, and the derivatives can be obtained analytically or with automatic differentiation [18, 19, 1, 20]. In case the function is nondifferentiable, so far, the most widespread approach is to linearize the problem such that it can still be solved with a gradient-based optimizer, in particular using the adjoint method [21, 11, 10, 22, 23, 24]. Although relatively successful, this approach comes with certain limitations, as highlighted for example in [9], where the authors use the adjoint formulation of the topology optimization problem to postpone fracture. The presented derivation of the adjoint assumes that the objective function is smooth, however, this is not the case for fracture, where even the smallest changes in the design can have a critical impact on the performance (specifically, these changes can completely change the crack trajectory). As highlighted by the authors, the effects cannot be easily remedied by fixing the step size of the descent algorithm or by regularization. In such a case, when the objective function is not well behaved, discontinuous, with multiple local minima, and difficult to regularize, gradient-free methods could be a promising alternative [13, 9, 16, 12].
Several gradient-free algorithms were explored in topology optimization, including Genetic Algorithms [25, 26, 27], Artificial Immune Algorithms [28], Ant Colonies [29, 30], Particle Swarms [31, 32], Simulated Annealing [33], Harmony Search [34], Differential Evolution [35] among others [13, 12]. These approaches, however, received strong criticism in the community [13] primarily due to their inferior efficiency, as opposed to gradient-based methods. Gradient-free algorithms suffer from the curse of dimensionality [13, 14], which means that the cost of optimization increases approximately exponentially with the number of problem dimensions. In standard topology optimization approaches [18], where the design variables correspond to element-wise densities, the dimensionality of even small ’toy-examples’ is in practice too large for gradient-free algorithms, leading to prohibitive costs (in order of 10,000 - 100,000 objective function evaluations [32], as opposed to the order of 10 – 100 evaluations for gradient-based methods [13]).
@@ 2.2 Latent-space optimization
The curse of dimensionality of gradient-free optimizers can be addressed by reparameterizing the optimization problem into a lower-dimensional latent space [12, 36, 37, 38]. This dimensionality reduction can be well accommodated using generative deep learning models, such as generative adversarial networks (GAN) [39], variational autoencoders (VAE) [40], and diffusion models [41, 42]. In the context of latent space optimization, variational autoencoders are a particularly common choice [43, 44].
Optimizing in the latent space of a VAE was demonstrated in the context of topology optimization [45, 46, 17, 47]. Sato et al. [45] and Gladstone et al. [46] explored latent optimization with a gradient-based approach using a surrogate model – a neural network trained to predict the property of interest from the latent representation. The method takes advantage of the differentiability of the surrogate – as the property of interest can be minimized by backpropagating through the surrogate. In both cases, the results demonstrate that the VAE is capable of representing new designs, that outperform those seen in the training. The disadvantage of this approach, however, is the need for data, required for training the surrogate.
Guo et al. [17] explore topology optimization in the latent space of a VAE with different optimization algorithms, including a gradient-free genetic algorithm. Unlike other works, the training of the VAE was augmented with an additional style-transfer procedure, to improve the quality of the generated designs. The training dataset comprised designs optimized for thermal compliance using the SIMP (solid isotropic material with penalization) method [18], but the model was then used to solve a different problem - a multi-objective optimization, minimizing the maximum temperature and maximizing the power density. Similar to other works [45, 46], the method was able to synthesize novel designs, different from those seen in the training dataset. Most importantly, the results demonstrate that the pre-trained model can be successfully deployed on a different optimization problem, given the designs have similar features. Nevertheless, in the presented experiments the optimization process with the gradient-free method (Genetic Algorithm) required an order of 30,000 FEM evaluations, indicating relatively limited improvement as compared to the gradient-free optimization without reparameterization into the latent space [13].
@@ 2.3 Machine learning in topology optimization
In the context of topology optimization, generative machine learning models have been primarily used to map boundary conditions to final designs [48, 6, 49]. Different generative models were explored in such a setup, including (GANs) [50, 51] and diffusion models [52]. While demonstrating competitive performance, these approaches are often criticized for their limited generalization [48, 6], as they remain applicable only to the problem on which they were trained. In order to apply such a model to a new type of design problem (e.g. different objective, different physics), the model needs to be retrained on a dataset corresponding to the new problem. Generating the dataset, however, requires the use of conventional optimization methods and simulations.
Another relevant stream of work is neural reparameterization [19, 53, 20, 54, 55, 6, 56, 57, 58, 59], where the neural network is used to implicitly represent the designed structure. Instead of optimizing the element-wise density values, the optimizer adjusts the trainable parameters of a neural network which outputs the density field. This approach proved beneficial in gradient-based approaches, where the weights can be adjusted by back-propagating the gradients of the density field (design sensitivities) [19]. Neural reparameterization was not explored in a gradient-free optimization scenario. One reason is the relatively high number of parameters of the neural networks, ranging from 100s [55] to 100,000s [60], making most network architectures unfeasible to optimize with gradient-free algorithms.
@@ 3 Methods
We propose a latent space optimization approach following two steps, as shown on the schematic in Figure 1. First, we train a generative model to reparameterize the topology designs into a lower-dimensional latent representation. Then, in the second step, the latent space is explored by the gradient-free optimizer, which attempts to find a design minimizing the objective. The optimizer controls the latent vector which parameterizes the design. This vector is decoded into a physical design using the pre-trained generative model, and the design is evaluated using the objective function (e.g. compliance). During one iteration of optimization, the optimizer generates a population of trial latent vectors, which are subsequently evaluated, and based on their associated objective values, the optimizer updates the population according to its internal update rules, converging to a better solution.
Figure 1: Schematics of our framework. a) Schematic of a variational autoencoder (VAE). In the first step of our method, we train the generative model – a VAE to reparameterize topology designs into latent space. Here $\mu$ and $\sigma$ are mean and variance vectors in latent space, $\epsilon$ is the noise vector sampled from a multivariate Gaussian, and $z$ is a latent variable. b) Schematic of latent space optimization process with gradient-free optimizer. In the second step of our method, we use a trained generative model, to optimize the designs using latent space representation of the generative model. Note that the latent vector $z$ is no longer stochastic, and it is fully controlled and updated ($z_{new}$) by the optimizer, based on the objective values obtained from the simulator.
The key assumption of our approach is that the generative model learns compact representations of features occurring in structural designs, that can be shared between designs optimized for different objectives (e.g. compliance, minimum mass, or compliant mechanism design). Evidence of this can be found in [9] or [15], by comparing designs optimized for fracture with those optimized for standard compliance. This observation can be verified using quantitative feature similarity metrics such as Learned Perceptual Image Patch similarity [61] (for details, see the appendix). Based on that, we assume that the generator trained to capture visual features can be deployed to optimize different objectives without the need for retraining and is not strictly limited to the problems used for generating the training dataset. This is a crucial advantage over state-of-the-art approaches exploring generative models for topology optimization.
@@ 3.1 Generative Model
We propose to use a new VAE variant called Latent Bernoulli Autoencoder (LBAE) that uses the Bernoulli distribution for modeling the latent space [62]. Although this architecture was proposed in a different context, it has been shown to achieve superior performance compared to conventional VAE architectures in terms of the quality of the generated samples (FID score) on a number of image generation tasks [62]. We conjectured that using a Bernoulli distribution for the latent space (zeros and ones), instead of a continuous distribution such as a Gaussian, is particularly adequate for the case of topology optimization where the aim is to decide whether each discrete element should contain material or not. We carefully compare the LBAE to a conventional VAE architecture and search for the best hyperparameters for a large class of problems, in an attempt to consider each method at their best performance [17]. Furthermore, we include gradient-based results and different gradient-free optimizers to establish clear performance metrics and clarify the usefulness of the proposed method.
@@
Conventionally, VAE provides a generative model that captures the distribution of data $\mathcal{D}=\left\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},...,\mathbf{x}^{(N)}\right\}$ assuming a joint distribution $p(\mathbf{x},\mathbf{z})$ using some hidden (latent) variable $\mathbf{z}$, such that given $\mathbf{z}$, it can generate samples of $\mathbf{x}$. In practice, this is achieved by using two neural networks: the encoder, which approximates the posterior $p(\mathbf{z}|\mathbf{x})\approx q_{\phi}(\mathbf{z}|\mathbf{x})$, (since in general $p(\mathbf{z}|\mathbf{x})$ is intractable), and the decoder, which captures $p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})$. The two networks are trained simultaneously by maximizing the evidence lower bound, as shown in Equation (1).
$L_{\mathbf{\theta},\mathbf{\phi}}=\mathbb{E}_{q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\frac{p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})p(\mathbf{z})}{q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})}\right]$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\left[\log p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})\right]-D_{KL}\left(q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})||p(\mathbf{z})\right)$ (1)
@@
The first term in Equation (1) corresponds to the reconstruction loss, and the second term corresponds to the Kullback-Leibler divergence, which regularizes the latent variable and aims to drive the parameterized distribution towards the assumed prior. In a conventional VAE, the prior $p(\mathbf{z})$ is assumed to be a multivariate Gaussian, while in LBAE the latent space is assumed to be a multivariate Bernoulli distribution. Unlike in VAE, however, in LBAE the Bernoulli distribution of the latent variable is imposed via network architecture and its specific sampling routine, rather than by an explicit regularization term (for details see [62]). Therefore, the second term of Equation (1) is not present in the LBAE loss.
@@
The training dataset $x\sim X$ in our case corresponds to images representing different topology designs. Once trained, the autoencoder is then used to generate new designs by sampling $z$ from the latent space. In this way, the designs can be represented using a lower-dimensional latent vector.
@@
The LBAE architecture that we use is based on the model proposed by Fajtl et al. [62], adapted to our dataset by adjusting the number of layers and convolutional filters. According to the authors [62], modeling latent space with Bernoulli distribution produces sharper images, compared to traditional VAE, and allows for smoother interpolation in the latent space. We hypothesize that these properties are crucial for optimization, allowing for easier traversing of the loss landscape by the optimizer. Furthermore, unlike in the baseline VAE, LBAE uses residual layers (for details, see Appendix E, Tables 2 and 3). Combined with the Bernoulli distribution of the latent variable, we observed that this architecture is indeed capable of producing output with finer detail and reduced blurriness, addressing one of the main shortcomings of VAE [6]. The baseline VAE architecture used in our case is a modified model of Larsen et al. [63]. The inputs and outputs are single-channel, grayscale representations of the designs of resolution 64x64.
For both LBAE and VAE, one of the most important parameters is the dimensionality of the latent space, which balances the training and the optimization performance. While increasing the latent space dimensionality makes the training easier, as the compression of information is not as severe, the optimization becomes less efficient due to the curse of dimensionality [13, 14]. To find a balanced latent space dimensionality, we tested several configurations (for details, see Appendix D). In the case of VAE, we did not observe a significant difference between the dimensionality 32 and 64, instead, we noticed that the performance was limited by insufficient resolution of finer details in the output (caused by the blurriness intrinsic to the VAE). These limiting factors, however, were to a large extent alleviated in LBAE, where the trade-off due to the latent space dimensionality was much more clear (for details, see the appendix). We found that for LBAE the dimensionality 256 offered the best balance.
@@
We use two different datasets of topology designs at train and test time, respectively. The training dataset was generated based on a modified approach of [64] and [52], resulting in 48,872 designs. The boundary conditions were generated as randomly drawn point loads and point supports, using Poisson’s distribution specified in the same way as in [64]. Additionally, in our dataset, we randomly applied a square mask to restrict part of the design domain from applying material (e.g., for an L-shaped beam, the mask would be applied in the corner). The mask was applied with 25% probability, and the position of the mask was randomly assigned along one of the edges of the design domain. Volume fractions were drawn using a uniform distribution with a range of [0.12, 0.5]. This dataset was split into two subsets: for training and validation, using a standard 80:20 split.
The testing dataset was used only for monitoring the testing loss, which aims to verify the generalization capability of the model. This dataset was generated using the problems defined in [19]. The boundary conditions were generated using one of 28 classes of parameterized problems, with the parameters drawn at random from a uniform distribution, which resulted in over 4000 instances of testing problems. Unlike the training problem set, the testing problem classes include distributed loads and supports; therefore, they are expected to provide a good insight into the generalization of the model and can be considered out-of-distribution.
The designs in both datasets were obtained using a gradient-based MMA (Method of Moving Asymptotes) [65] optimizer with the default hyperparameter settings. The final designs were represented as grayscale images of fixed resolution (64x64) with pixel values representing the volume fraction of the material. For each problem, the optimization with MMA [65] was run for 200 iterations. The simulations were set up following the approach of the 88-line code [18], with $E_{0}=1.0$, $E_{min}=1e-9$, $\nu=0.3$ and a standard cone filter with a radius of 2 pixels to regularize the density field.